OpenStudy (njabulo):

find the derivative of y = (1/ log2^x - arcsinx)^2

OpenStudy (anonymous):

brace yourself 2 (-1/sqrt(1-x^2)-2/(x log^3(x))) (1/(log^2(x))-sin^(-1)(x))

OpenStudy (anonymous):

-2$\big(\dfrac{1}{x\log a}-\sin^{-1}x\big)\big(\dfrac{1}{x^2\log a}+\dfrac{1}{\sqrt{1-x^2}}\big)$